# MMA-UNet: A Multi-Modal Asymmetric UNet Architecture for Infrared and Visible Image Fusion

### v1

@@ 1. Introduction
  • We proposed a multi-modal asymmetric UNet (MMA-UNet) architecture for IVIF that uses a simple and efficient method for fusing multi-modal features when compared with the existing complex paradigms.
@@
1.1. Motivation
The existing IVIF algorithms (Li et al., 2023c; Xu et al., 2023; Liu et al., 2022; Huang et al., 2022; Liu et al., 2023a; Zhang et al., 2020) primarily focus on efficiently learning the multi-modal feature interactions while ignoring the inconsistency of the spatial information distribution in multi-modal image. Therefore, we rethought the multi-modal fusion paradigm. Firstly, the two unique UNets were trained separately on IR and VI dataset, represented by IR-UNet and VI-UNet. We then calculated the Centered Kernel Alignment (CKA) (Nguyen et al., 2020) similarity of the features extracted by the encoders of the two UNets. As shown in fig. 1(a) and (b), note that the total number of layers is much greater than the stated depth of the UNet. The latter only accounts for the convolutional layers in the network, but we include all intermediate representations. Specifically, for the UNet framework, the first convolution block contains 7 layers of feature intermediate representation, and the second to fifth convolution blocks contain 11 layers of feature intermediate representation. Neural networks tend to initially learn shallow features, and as the network depth increases, they gradually dive into the deeper semantic spaces for feature learning. The shallow features are mostly similar. However, as the network depth increases, the differences between the deep and shallow features increase. IR captures shallow features in the first 22 layers since they share a greater similarity with the surrounding features, as shown in fig. 1(a). When the network depth exceeds 22 layers, the extracted features vary significantly from the surroundings, manifested as deeper colors in the fig. 1. Conversely, the feature extraction network of the VI begins to show its uniqueness from the 12th layer, and the difference in the subsequent features continues to increase. Therefore, we observed that VI can reach deeper semantic spaces faster than IR under the same architecture, and VI can extract shallow information 10 layers faster than IR in our network.
@@ 2.1. Infrared and Visible Image Fusion Methods
The existing deep learning-based IVIF algorithms are primarily divided into two categories: generative and encoding models. Generative models are primarily based on the generative adversarial network (GAN) (Le et al., 2022; Ma et al., 2020b, 2019b; Rao et al., 2023a; Ma et al., 2020a) and diffusion (Zhao et al., 2023c; Yue et al., 2023). They aimed to learn data distribution from the latent space and simulate the distribution of the target data by generating the data. For example, Liu et al. (Liu et al., 2022) proposed a target-aware dual adversarial learning strategy to generate fusion results that are beneficial for downstream tasks. They designed the target and detailed discriminators to learn the target features from IR and VI, respectively. To address the issues of weak training stability and lack of interpretability in GAN-based models, Zhao et al. (Zhao et al., 2023c) introduced a denoising diffusion model for IVIF, defining the fusion task as an unconditional generation sub-problem and a maximum likelihood sub-problem. Conversely, encoding models do not generate new data samples and instead, extract key features from the original input data, mapping the original data to a more compact representation space (Ma et al., 2022; Tang et al., 2022c; Rao et al., 2023b; Liang et al., 2022; Liu et al., 2021; Chang et al., 2023; Tang et al., 2022b). For instance, Ma et al. (Ma et al., 2022) introduced Swin Transformer to the image fusion, incorporating an attention-guided cross-domain module. This design effectively integrates multi-modality complementary information and global interaction information. To enhance the interpretability of feature extraction, Li et al. (Li et al., 2023c) proposed an image fusion framework based on representation learning, that established a connection between the mathematical formulas and network architecture to enhance the interpretability of feature extraction. They utilized low-rank representation learning theory to establish an image decomposition model, effectively avoiding a time-consuming network design. Owing to the continuous improvement in image fusion frameworks, several researchers have begun to analyze the performance of fusion models in promoting downstream tasks. The many multitask joint learning methods have emerged (Liu et al., 2022; Tang et al., 2022c). For example, fusion and object detection (Liu et al., 2022), fusion and semantic segmentation (Tang et al., 2022c; Liu et al., 2023b), fusion and salient object detection (Wang et al., 2023b), and fusion and low-light enhancement (Tang et al., 2023b). They generally guide the learning of fusion networks through the performance feedback of the fusion results on downstream tasks. Additionally, several works proposed ”registration and fusion” models (Xu et al., 2023; Wang et al., 2022; Xie et al., 2023) due to the spatial deformation and misalignment in the imaging process of multi-modal sensors to effectively avoid the excessive reliance of algorithms on the registration data pairs.
@@ 5.1. Conclusion
In this study, we proposed an asymmetric UNet architecture for IVIF, providing a simple and efficient method for fusing multi-modal image features. We analyze the differences in the spatial information distribution between the IR and VI modalities, and duduce the conclusion that there are discrepancy in the speed of extracting deep semantic features from different modalities within the same framework. We designed a cross-scale fusion rule based on different layer counts to address this issue. Subsequently, we designed a guidance mechanism for training IR-UNet and observed that simple guidance using VI features enhanced the efficiency of the deep semantic feature extraction in IR. The experimental results demonstrate that MMA-UNet outperforms the existing mainstream architectures, achieving good performance in MMIF and downstream tasks.

### v2

@@ 1. Introduction
  • We proposed a multi-modal asymmetric UNet (MMA-UNet) architecture for IVIF that uses a simple and effective method for fusing multi-modal features when compared with the existing complex paradigms.
@@
Motivation. The existing IVIF algorithms (Li et al., 2023c; Xu et al., 2023; Liu et al., 2022; Huang et al., 2022; Liu et al., 2023a; Zhang et al., 2020) primarily focus on effectively learning the multi-modal feature interactions while ignoring the inconsistency of the spatial information distribution in multi-modal image. Therefore, we rethought the multi-modal fusion paradigm. Firstly, the two unique UNets were trained separately on IR and VI dataset, represented by IR-UNet and VI-UNet. We then calculated the Centered Kernel Alignment (CKA) (Nguyen et al., 2020) similarity of the features extracted by the encoders of the two UNets. As shown in fig. 1(a) and (b), note that the total number of layers is much greater than the stated depth of the UNet. The latter only accounts for the convolutional layers in the network, but we include all intermediate representations. Specifically, for the UNet framework, the first convolution block contains 7 layers of feature intermediate representation, and the second to fifth convolution blocks contain 11 layers of feature intermediate representation. Neural networks tend to initially learn shallow features, and as the network depth increases, they gradually dive into the deeper semantic spaces for feature learning. The shallow features are mostly similar. However, as the network depth increases, the differences between the deep and shallow features increase. IR captures shallow features in the first 22 layers since they share a greater similarity with the surrounding features, as shown in fig. 1(a). When the network depth exceeds 22 layers, the extracted features vary significantly from the surroundings, manifested as deeper colors in the fig. 1. Conversely, the feature extraction network of the VI begins to show its uniqueness from the 12th layer, and the difference in the subsequent features continues to increase. Therefore, we observed that VI can reach deeper semantic spaces faster than IR under the same architecture, and VI can extract shallow information 10 layers faster than IR in our network.
@@ 2. Related Work
Infrared and Visible Image Fusion Methods. The existing deep learning-based IVIF algorithms are primarily divided into two categories: generative and encoding models. Generative models are primarily based on the generative adversarial network (GAN) (Le et al., 2022; Ma et al., 2020b, 2019b; Rao et al., 2023a; Ma et al., 2020a) and diffusion (Zhao et al., 2023c; Yue et al., 2023). They aimed to learn data distribution from the latent space and simulate the distribution of the target data by generating the data. For example, Liu et al. (Liu et al., 2022) proposed a target-aware dual adversarial learning strategy to generate fusion results that are beneficial for downstream tasks. They designed the target and detailed discriminators to learn the target features from IR and VI, respectively. To address the issues of weak training stability and lack of interpretability in GAN-based models, Zhao et al. (Zhao et al., 2023c) introduced a denoising diffusion model for IVIF, defining the fusion task as an unconditional generation sub-problem and a maximum likelihood sub-problem. Conversely, encoding models do not generate new data samples and instead, extract key features from the original input data, mapping the original data to a more compact representation space (Ma et al., 2022; Tang et al., 2022c; Rao et al., 2023b; Liang et al., 2022; Liu et al., 2021; Chang et al., 2023; Tang et al., 2022b). For instance, Ma et al. (Ma et al., 2022) introduced Swin Transformer to the image fusion, incorporating an attention-guided cross-domain module. This design effectively integrates multi-modality complementary information and global interaction information. To enhance the interpretability of feature extraction, Li et al. (Li et al., 2023c) proposed an image fusion framework based on representation learning, that established a connection between the mathematical formulas and network architecture to enhance the interpretability of feature extraction. They utilized low-rank representation learning theory to establish an image decomposition model, effectively avoiding a time-consuming network design. Owing to the continuous improvement in image fusion frameworks, several researchers have begun to analyze the performance of fusion models in promoting downstream tasks. The many multitask joint learning methods have emerged (Liu et al., 2022; Tang et al., 2022c). For example, fusion and object detection (Liu et al., 2022), fusion and semantic segmentation (Tang et al., 2022c; Liu et al., 2023b), fusion and salient object detection (Wang et al., 2023b), and fusion and low-light enhancement (Tang et al., 2023b). They generally guide the learning of fusion networks through the performance feedback of the fusion results on downstream tasks. Additionally, several works proposed ”registration and fusion” models (Xu et al., 2023; Wang et al., 2022; Xie et al., 2023) due to the spatial deformation and misalignment in the imaging process of multi-modal sensors to effectively avoid the excessive reliance of algorithms on the registration data pairs.
@@ 5.1. Conclusion
In this study, we proposed an asymmetric UNet architecture for IVIF, providing a simple and effective method for fusing multi-modal image features. We analyze the differences in the spatial information distribution between the IR and VI modalities, and duduce the conclusion that there are discrepancy in the speed of extracting deep semantic features from different modalities within the same framework. We designed a cross-scale fusion rule based on different layer counts to address this issue. Subsequently, we designed a guidance mechanism for training IR-UNet and observed that simple guidance using VI features enhanced the efficiency of the deep semantic feature extraction in IR. The experimental results demonstrate that MMA-UNet outperforms the existing mainstream architectures, achieving good performance in MMIF and downstream tasks.